# algorithmic

### v1

@@ 4 Aggregation-based GNNs
where $\mathbf{h}^{(0)}_{i}$ is initialized with the feature vector of node $i$, $\mathrm{AGG}^{(l)}(\cdot)$ is instantiated as a set-based pooling operation such as MAX, MEAN [19], or attention-based SUM [61, 36], $\mathbf{W}^{(l)}$ is a weight matrix for the $l$-th GNN layer, which is shared across all nodes and used for representation transformation (i.e., if $\hat{\mathbf{h}}^{(l)}_{v}\in\mathbb{R}^{f},\mathbf{W}^{(l)}\in\mathbb{R}^{f\times f^{\prime}}$, then $\mathbf{h}^{(l)}_{v}\in\mathbb{R}^{f^{\prime}}$). For simplicity, we omit the residual connections, activation functions, etc. In this paper, we use the term GNNs to refer to such aggregation-based GNNs unless otherwise stated.
@@
where $a_{i,j}^{(l)}$ is the weight for the message (i.e., $\mathbf{h}^{(l-1)}_{j}$) from node $j$ to $i$. For MEAN-pooling in GNNs like GCN [19], it can be $a_{i,j}^{(l)}=1,\forall\mathcal{E}_{i,j}\in\mathbb{E}$. For attention-based GNNs like GAT [61], $a_{i,j}^{(l)}$ is an attention coefficient that is computed dynamically based on $\mathbf{h}^{(l-1)}_{i}$ and $\mathbf{h}^{(l-1)}_{j}$.
@@
Analysis. As shown in Eq. 1, GNNs update the representation $\mathbf{h}^{(l)}_{i}$ by aggregating the representations of node $i$ and its neighbors. In this process, the number of neighbors of node $i$ can hardly be encoded into $\mathbf{h}^{(l)}_{i}$. This is due to the inherent nature of neighborhood aggregation scheme in GNNs, i.e., the $\mathrm{AGG}^{(l)}(\cdot)$ in Eq. 1 is set-based pooling, which is originally designed to handle irregular sizes of neighborhood sets of different nodes in a graph. For example, if the aggregation is MEAN pooling, then the set of node-wise representations (i.e., $\{\mathbf{h}^{(l-1)}_{j}~{}|~{}\forall j\in\Gamma_{i}\cup\{i\}\}$ in Eq. 1) will be averaged and the result could hardly contain information about the size of that set. Note that attention-based pooling also cannot address this inherent issue of the neighborhood aggregation scheme. As shown in Eq. 2, attention-based GNNs [61] essentially replace the original edge weight with attention weight. Despite this modification, the set of representations is weighted averaged, and consequently, the resulting representation still lacks information related to the size of the neighborhood set.
Remark 2 shows that the neighborhood aggregation algorithm of GNNs inherently cannot effectively learn information about the number of neighbors of each node. Essentially, we can address this issue by, for example, adding the node degree as a feature to each node. We note that several previous works have pointed out this [67, 76]. We present it using Remark 2 for better presenting our following analysis.
5 Can GNNs Completely Learn NCN-dependent Structural Information?
5.1 Analytical Study
What can we do when applying the node-wise representations learned by GNNs to downstream graph tasks that involve multiple nodes, such as link prediction or graph classification? A general way is to combine the representations of the involved nodes into one representation and pass it into the subsequent model components [67, 63, 36, 27]. For such a combination, we have the following insight:
@@ 5.1 Analytical Study
Analysis. According to Remark 2, due to the inherent nature of the neighborhood aggregation algorithm, node-wise representations learned by GNNs cannot effectively capture information related to the number of neighbors of each node (i.e., the size of its neighborhood set), much less to the number of common neighbors between two nodes. The operation of combining representations of two or more nodes also cannot effectively extract NCN-dependent structural information. For example, we can combine the representations of two nodes by concatenation, Hadamard production [63, 27], and combine more nodes’ representations by MEAN pooling [67], Sort pooling [75], etc. These combination operations on node-wise representations learned by GNNs are unlikely to contain the behavior of counting the common neighbors between two nodes, thereby falling short in extracting NCN-dependent structural information.
GNNs might learn little structural information related to the number of common neighbors. However, the neighborhood aggregation algorithm of GNNs learns node-wise representations by passing the messages of neighboring nodes of each node to that node and set-based aggregates them [19, 61, 76, 36]. Such set-based aggregation operation inherently washes out the information related to the number of nodes in the set, including the number of common nodes between two nodes. For instance, as shown in Fig. 1, the nodes $1$ and $4$ have common neighboring nodes $5,6$. In GNN learning based on Eq. 1, the node $1$ will receive the messages from $2,3,5,6$, where the number of common neighbors (i.e., $\mathrm{CN}_{1,4}=2$) can hardly be captured in the aggregation of five representations (i.e., representations of nodes $1,2,3,4,5$). Note that in this example, attention-based aggregation also cannot effectively learn $\mathrm{CN}_{1,4}=2$ from the aggregation of five representations. The reason is the same as the analysis of Remark 2. In fact, it is difficult to interpret what information the GNN has learned in a rigorous mathematical format. Nevertheless, we could say that GNNs cannot completely capture NCN-dependent structural information.
@@
5.2 Empirical Study
5.2.1 Experimental Design
If Remark 3 holds, we expect that properly integrating NCN-dependent heuristics into a GNN could improve the link prediction performance. To this end, we design our experiments as detailed in Algorithm 2. As shown, given two nodes $v,u$, the link prediction is performed by combining the two nodes’ representations from the last GNN layer into a pair-specific link representation $\mathbf{h}_{vu}$, then concatenating it with heuristic encodings, and lastly passing the concatenation into a predictor like MLP. During the training stage, the node pair $(v,u)$ can be a positive or negative link sample, where a negative sample can be two distant nodes that are not connected to each other.
In Algorithm 2, $\mathbf{h}^{(0)}_{i}$ can be initialized using feature vector $\mathbf{x}_{i}\in\mathbf{X}$, node embedding $\mathbf{e}_{i}\in\mathbf{E}$ or the concatenation of both $\mathbf{x}_{i}$ and $\mathbf{e}_{i}$. Some may refer to node embedding as an intermediate representation of a node in GNNs. In this work, we clearly distinguish node embeddings from node representations. We consider node embedding as a type of node-wise input feature. The embedding of a node can be viewed as encoding a unique node id into a trainable embedding vector, which is like encoding a unique word id into the word embedding in natural language processing [45]. Note that we can encode any feature into a trainable embedding vector (e.g., encoding node degree to an embedding). The main difference between node embeddings and the embeddings of other node features is that each node embedding vector is unique to that node. For example, an embedding of a node degree is not unique to that node (different nodes could have the same node degree).
@@ 5.2.1 Experimental Design
5.2.2 Datasets
Table 2: Statistics of OGB link prediction datasets used in our experiments.
@@ 5.2.2 Datasets
ogbl-citation2 $2,927,963$ $30,561,187$ $21$
All of our experiments are conducted on four OGB link prediction datasets: ogbl-collab, ogbl-citation2, ogbl-ppa, and ogbl-ddi [21]. The statistics of datasets are summarized in Table 2. All these datasets are constructed based on real-world data, covering diverse realistic applications and spanning different scales (4K - 3M nodes). OGB provides an official evaluation protocol. We completely follow it in the data splits and evaluation metrics (i.e., Hits@$50$, MRR, Hits@$100$, and Hits@$20$ on ogbl-collab, ogbl-citation2, ogbl-ppa and ogbl-ddi, respectively). We report the result on the test set, with mean and standard deviation computed across 10 trials.
Figure 2: The results of Algorithm 2 on four OGB link prediction datasets, using heuristic encoding (HE) only, node features (X) only, node embeddings (NE) only, or their combinations. The data splits and evaluation metrics follow OGB official evaluation protocol [21].
5.2.3 Implementation Details
We implement our Algorithm 2 based on PyTorch and PyTorch Geometric [13]. All embedding vectors are initialized by following the methods in [16, 20]. The model is trained with Adam optimizer [28]. The learning rate is decayed using the ExponentialLR method [38]. We conduct experiments on ogbl-ddi and ogbl-collab on a Linux machine with 192G RAM, and NVIDIA Quadro P6000 (24G), and on ogbl-ppa and ogbl-citation2 on a machine with 512G RAM and NVIDIA A100 (40G). Table 3 lists the configurations of the Algorithm 2 for the best performance. We provide our code for reproducing the results at https://github.com/astroming/GNNHE.
Table 3: Configurations of the Algorithm 2 for the best performances.
@@ 5.2.4 Experimental Results
6 GNNs with Node Embedding in Link Prediction
@@
Analysis. In Algorithm 2, the parameters optimized by the link samples could include model weights, trainable node embeddings, and other embedding weights. The key difference between node embedding weights and other learnable weights is that the former is unique to each node but the latter is shared across multiple nodes (e.g., the GNN weight matrix $\mathbf{W}^{(l)}$ in Eq. 1 is shared across all nodes). The unique nature of node embeddings can bring benefits. As shown in Algorithm 2, when the model training is supervised by a link sample $(v,u)$, for a GNN using node embeddings, the loss calculated based on $(y_{v,u},\hat{y}_{v,u})$ would be used to optimize the node embeddings of nodes $v,u$ and their neighboring nodes (i.e., the nodes involved in calculating $\mathbf{h}^{(L)}_{v},\mathbf{h}^{(L)}_{u}$). The link state of $(v,u)$ could be encoded into the node embeddings of these nodes, which would enable those node embeddings to remember the relationships between corresponding nodes. After being trained with sufficient positive and negative link samples, the node embedding of each node could know which nodes (through their node embeddings) in the graph are more likely to be or not to be connected to that node.
@@
Finding 5 indicates that the graph degree significantly influences the effectiveness of trainable node embeddings in GNN-based link prediction models. Interestingly, prior studies [54] have also highlighted the sensitivity of heuristic methods to the graph degree. This underscores the necessity of considering the graph degree when selecting link prediction methods, as their efficacy may vary depending on it. Investigating the influence of different graph degrees on link prediction methods represents a compelling direction for further research.
7 GNNs in Learning non-NCN-dependent Structural Information
In Section 5, we present that GNNs cannot completely learn the structural information related to the number of common neighbors between two target nodes. When it comes to the question of what non-NCN-dependent pair-specific structural information can be learned via GNNs, it poses a significant challenge. This is due to the potential presence of diverse types of non-NCN-dependent information. Unlike NCN-dependent information directly related to the number of common neighbors, non-NCN-dependent information tends to be abstract and difficult to express in rigorous mathematical terms. For example, our review in Section 3.2 identifies only three non-NCN-dependent link heuristics. In this work, we leave the exploration of this question as a future research endeavor. Nevertheless, by comparing Algorithm 1 and Algorithm 2, we have the following insight.
@@ 7 GNNs in Learning non-NCN-dependent Structural Information
8 Limitation Analysis of Existing Methods
@@
8.1 A Survey of Link Prediction Methods
8.1.1 Heuristic Methods
@@
As outlined in Section 3, traditional link heuristics are usually defined based on the number of common neighbors or paths between two nodes [43]. Their effectiveness in link prediction has been confirmed in real-world tasks [43, 31]. However, many link heuristics are designed for specific graph applications and their performance may vary on different graphs [31]. Also, the expressiveness of these methods is limited compared to graph representation learning [74].
8.1.2 Graph Neural Networks
@@
GNNs have proven their effectiveness in various graph applications [39, 22, 71, 27]. A number of GNN models have been proposed [29, 68, 83]. GCN  [29] learns node representations by summing the normalized representations from the first-order neighbors. GraphSAGE [19] samples and aggregates representations from local neighborhoods. GAT [61] introduces an attention-based GNN architecture. JKNet [68] adds a pooling layer following the last GNN layer and each GNN layer has a residual connection to this layer. Cluster-GCN [10] proposes an efficient algorithm for training deep GCN on large graphs. LRGA [48] incorporates a Low-Rank global attention module to GNNs. Several works such as Mixhop [1] and DEGNN [37] propose techniques to leverage high-hop neighbors. ID-GNN [72] embeds each node by considering its identity.These GNNs have demonstrated promising link prediction performance.
8.1.3 Non-GNN-based Node Embedding Methods
@@
A family of node embedding methods is those built with matrix factorization [30]. MF [44] is a pioneer work employing matrix factorization in link prediction. FSSDNMF [8] proposes a link prediction model based on non-negative matrix factorization. In general, such methods mainly rely on the adjacency matrix and tend to encounter scalability issues when employed on large graphs.
@@
Another family of node embedding methods is those based on relative distance encoding. The similarity of nodes in the embedding space reflects the semantic similarity of nodes in the graph [47]. Such methods would learn more similar embeddings for two close nodes than two distant nodes. Following word embedding [45], methods such as Deepwalk [47], Node2vec [17], and NodePiece [14] learn node embeddings by treating the nodes as words and treating the sequences of nodes generated based on links as sentences. UniNet [69] improves the efficiency of such methods using the Metropolis Hastings sampling technique [11]. Inspired by subword tokenization [53], NodePiece [14] explores parameter-efficient node embeddings.
8.1.4 SEAL-type Methods

### v2

@@ IV Aggregation-based GNNs
where $\mathbf{h}^{(0)}_{i}$ is initialized with the feature vector of node $i$, $\mathrm{AGG}^{(l)}(\cdot)$ is instantiated as a set-based pooling operation such as MAX, MEAN [sage2017inductive], or attention-based SUM [gat2018graph, li2024evaluating], $\mathbf{W}^{(l)}$ is a weight matrix for the $l$-th GNN layer, which is shared across all nodes and used for representation transformation (i.e., if $\hat{\mathbf{h}}^{(l)}_{v}\in\mathbb{R}^{f},\mathbf{W}^{(l)}\in\mathbb{R}^{f\times f^{\prime}}$, then $\mathbf{h}^{(l)}_{v}\in\mathbb{R}^{f^{\prime}}$). For simplicity, we omit the residual connections, activation functions, etc. In this paper, we use the term GNNs to refer to such aggregation-based GNNs unless otherwise stated.
@@
[…]
@@
Analysis. As shown in Eq. 1, GNNs update the representation $\mathbf{h}^{(l)}_{i}$ by aggregating the representations of node $i$ and its neighbors. In this process, the number of neighbors of node $i$ can hardly be encoded into $\mathbf{h}^{(l)}_{i}$. This is due to the nature of the neighborhood aggregation scheme in GNNs, i.e., the $\mathrm{AGG}^{(l)}(\cdot)$ in Eq. 1 is set-based pooling, which is originally designed to handle irregular sizes of neighborhood sets of different nodes in a graph. For example, if the aggregation is MEAN pooling, then the set of node-wise representations (i.e., $\{\mathbf{h}^{(l-1)}_{j}~|~\forall j\in\Gamma_{i}\cup\{i\}\}$ in Eq. 1) will be averaged and the result could hardly contain information about the size of that set. Note that attention-based pooling also cannot address this inherent issue of the neighborhood aggregation scheme. As shown in Eq. 2, attention-based GNNs [gat2018graph] essentially replace the original edge weight with attention weight. Despite this modification, the set of representations is weighted averaged, and consequently, the resulting representation still lacks information related to the size of the neighborhood set.
Remark 2 shows that the neighborhood aggregation algorithm of GNNs cannot effectively learn information about the number of neighbors of each node. Essentially, we can address this issue by, for example, adding the node degree as a feature to each node. We note that several previous works have pointed out this [xu2018powerful, zhang2021labeling]. We present it using Remark 2 for better presenting our following analysis.
V Can GNNs Completely Learn NCN-dependent Structural Information?
V-A Analytical Study
What can we do when applying the node-wise representations learned by GNNs to downstream graph tasks that involve multiple nodes, such as link prediction or graph classification? A general way is to combine the representations of the involved nodes into one representation and pass it into the subsequent model components [xu2018powerful, wang2021pairwise, li2024evaluating, khemani2024review]. For such a combination, we have the following insight:
@@ V-A Analytical Study
Analysis. According to Remark 2, due to the nature of the neighborhood aggregation algorithm, node-wise representations learned by GNNs cannot effectively capture information related to the number of neighbors of each node (i.e., the size of its neighborhood set), much less to the number of common neighbors between two nodes. The operation of combining representations of two or more nodes also cannot effectively extract NCN-dependent structural information. For example, we can combine the representations of two nodes by concatenation, Hadamard production [wang2021pairwise, khemani2024review], and combine more nodes’ representations by MEAN pooling [xu2018powerful], Sort pooling [zhang2018end], etc. These combination operations on node-wise representations learned by GNNs are unlikely to contain the behavior of counting the common neighbors between two nodes, thereby falling short in extracting NCN-dependent structural information.
GNNs might learn little structural information related to the number of common neighbors. However, the neighborhood aggregation algorithm of GNNs learns node-wise representations by passing the messages of neighboring nodes of each node to that node and set-based aggregates them [sage2017inductive, gat2018graph, zhang2021labeling, li2024evaluating]. Such set-based aggregation operation inherently washes out the information related to the number of nodes in the set, including the number of common nodes between two nodes. For instance, as shown in Fig. 1, the nodes $1$ and $4$ have common neighboring nodes $5,6$. In GNN learning based on Eq. 1, the node $1$ will receive the messages from $2,3,5,6$, where the number of common neighbors (i.e., $\mathrm{CN}_{1,4}=2$) can hardly be captured in the aggregation of five representations (i.e., representations of nodes $1,2,3,4,5$). Note that in this example, attention-based aggregation also cannot effectively learn $\mathrm{CN}_{1,4}=2$ from the aggregation of five representations. The reason is the same as the analysis of Remark 2. In fact, it is difficult to interpret what information the GNN has learned in a rigorous mathematical format (an acknowledged limitation of our study (see Section X)). Nevertheless, we could say that GNNs cannot completely capture NCN-dependent structural information.
@@
V-B Empirical Study
V-B1 Experimental Design
If Remark 3 holds, we expect that properly integrating NCN-dependent heuristics with a GNN could improve the link prediction performance. To this end, we design our experiments as detailed in Algorithm 2. As shown, given two nodes $v,u$, the link prediction is performed by combining the two nodes’ representations from the last GNN layer into a pair-specific link representation $\mathbf{h}_{vu}$, then concatenating it with heuristic encodings, and lastly passing the concatenation into a predictor like MLP. During the training stage, the node pair $(v,u)$ can be a positive or negative link sample, where a negative sample can be two distant nodes that are not connected to each other.
In Algorithm 2, $\mathbf{h}^{(0)}_{i}$ can be initialized using feature vector $\mathbf{x}_{i}\in\mathbf{X}$, node embedding $\mathbf{e}_{i}\in\mathbf{E}$ or the concatenation of both $\mathbf{x}_{i}$ and $\mathbf{e}_{i}$. Some may refer to node embedding as an intermediate representation of a node in GNNs. In this work, we clearly distinguish node embeddings from node representations. We consider node embedding as a type of node-wise input feature. The embedding of a node can be viewed as encoding a unique node id into a trainable embedding vector, which is like encoding a unique word id into the word embedding in natural language processing [mikolov2013distributed]. Note that we can encode any feature into a trainable embedding vector (e.g., encoding node degree to an embedding). The main difference between node embeddings and the embeddings of other node features is that each node embedding vector is unique to that node. For example, an embedding of a node degree is not unique to that node (different nodes could have the same node degree).
@@ V-B1 Experimental Design
V-B2 Datasets
TABLE III: Statistics of OGB link prediction datasets used in our experiments.
@@ V-B2 Datasets
ogbl-citation2 $2,927,963$ $30,561,187$ 2121
All of our experiments are conducted on four OGB link prediction datasets: ogbl-collab, ogbl-citation2, ogbl-ppa, and ogbl-ddi [hu2020open]. The statistics of datasets are summarized in Table III. All these datasets are constructed based on real-world data, covering diverse realistic applications and spanning different scales (4K - 3M nodes). OGB provides an official evaluation protocol. We completely follow it in the data splits and evaluation metrics (i.e., Hits@$50$, MRR, Hits@$100$, and Hits@$20$ on ogbl-collab, ogbl-citation2, ogbl-ppa and ogbl-ddi, respectively). We report the result on the test set, with mean and standard deviation computed across 10 trials.
Figure 2: The results of Algorithm 2 on four OGB link prediction datasets, using heuristic encoding (HE) only, node features (X) only, node embeddings (NE) only, or their combinations. The data splits and evaluation metrics follow OGB official evaluation protocol [hu2020open].
V-B3 Implementation Details
We implement our Algorithm 2 based on PyTorch and PyTorch Geometric [FeyLenssen2019]. All embedding vectors are initialized by following the methods in [glorot2010understanding, he2015delving]. The model is trained with Adam optimizer [kingma2014adam]. The learning rate is decayed using the ExponentialLR method [Li2020An]. We conduct experiments on ogbl-ddi and ogbl-collab on a Linux machine with 192G RAM, and NVIDIA Quadro P6000 (24G), and on ogbl-ppa and ogbl-citation2 on a machine with 512G RAM and NVIDIA A100 (40G). Table IV lists the configurations of the Algorithm 2 for the best performance. We provide our code for reproducing the results at https://github.com/astroming/GNNHE.
TABLE IV: Configurations of the Algorithm 2 for the best performances.
@@ V-B4 Experimental Results
VI GNNs with Node Embedding in Link Prediction
@@
[…]
@@
Finding 5 indicates that the graph degree significantly influences the effectiveness of trainable node embeddings in GNN-based link prediction models. Interestingly, prior studies [shang2019link] have also highlighted the sensitivity of heuristic methods to the graph degree. This underscores the necessity of considering the graph degree when selecting link prediction methods, as their efficacy may vary depending on it. Investigating the influence of different graph degrees on link prediction methods represents a compelling direction for further research.
VII GNNs in Learning non-NCN-dependent Structural Information
In Section V, we present that GNNs cannot completely learn the structural information related to the number of common neighbors between two target nodes. When it comes to the question of what non-NCN-dependent pair-specific structural information can be learned via GNNs, it poses a significant challenge. This is due to the potential presence of diverse types of non-NCN-dependent information. Unlike NCN-dependent information directly related to the number of common neighbors, non-NCN-dependent information tends to be abstract and difficult to express in rigorous mathematical terms. For example, PA relies on the product of node degrees, reflecting a commonly observed phenomenon in social and web networks where newly added nodes are more likely to connect to nodes with high degrees. Our review in Section III-B highlights three widely used non-NCN-dependent link heuristics. In this work, we leave the exploration of this question as a future research endeavor. Nevertheless, by comparing Algorithm 1 and Algorithm 2, we have the following insight.
@@ VII GNNs in Learning non-NCN-dependent Structural Information
VIII Limitation Analysis of Existing Methods
@@
VIII-A A Survey of Link Prediction Methods
VIII-A1 Heuristic Methods
@@
As outlined in Section III, traditional link heuristics are usually defined based on the number of common neighbors or paths between two nodes [martinez2016survey]. Their effectiveness in link prediction has been confirmed in real-world tasks [martinez2016survey, kovacs2019network]. However, many link heuristics are designed for specific graph applications and their performance may vary on different graphs [kovacs2019network]. Also, the expressiveness of these methods is limited compared to graph representation learning methods [zhang2018link].
VIII-A2 Graph Neural Networks
@@
GNNs have proven their effectiveness in various graph applications [liang2021failure, huang2020skipgnn, ying2021transformers, khemani2024review]. A number of GNN models have been proposed [kipf2017semi, xu2018representation, zou2023simple]. GCN  [kipf2017semi] learns node representations by summing the normalized representations from the first-order neighbors. GraphSAGE [sage2017inductive] samples and aggregates representations from local neighborhoods. GAT [gat2018graph] introduces an attention-based GNN architecture. JKNet [xu2018representation] adds a pooling layer following the last GNN layer and each GNN layer has a residual connection to this layer. Cluster-GCN [chiang2019cluster] proposes an efficient algorithm for training deep GCN on large graphs. LRGA [puny2020global] incorporates a Low-Rank global attention module to GNNs. Several works such as Mixhop [abu2019mixhop] and DEGNN [li2020distance] propose techniques to leverage high-hop neighbors. ID-GNN [you2021identity] embeds each node by considering its identity. These GNNs have demonstrated promising link prediction performance.
VIII-A3 Non-GNN-based Node Embedding Methods
@@
A family of node embedding methods is those built with matrix factorization [koren2009matrix]. MF [menon2011link] is a pioneer work employing matrix factorization in link prediction. FSSDNMF [chen2022link] proposes a link prediction model based on non-negative matrix factorization. In general, such methods mainly rely on the adjacency matrix and tend to encounter scalability issues when employed on large graphs.
@@
Another family of node embedding methods is those based on relative distance encoding. The similarity of nodes in the embedding space reflects the semantic similarity of nodes in the graph [perozzi2014deepwalk]. Such methods would learn more similar embeddings for two close nodes than two distant nodes. Following word embedding [mikolov2013distributed], methods such as Deepwalk [perozzi2014deepwalk], Node2vec [grover2016node2vec], and NodePiece [galkin2021nodepiece] learn node embeddings by treating the nodes as words and treating the sequences of nodes generated based on links as sentences. UniNet [yao2021uninet] improves the efficiency of such methods using the Metropolis Hastings sampling technique [chib1995understanding]. Inspired by subword tokenization [sennrich2016neural], NodePiece [galkin2021nodepiece] explores parameter-efficient node embeddings.
VIII-A4 SEAL-type Methods